Documenta Mathematica

Band 5, 2000

Atallah Affane Formules de Représentation Intégrale

pour les Domaines de Cartan  1–13

Karl-Heinz Hoffmann, Victor N. Starovoitov Zur Bewegung einer Kugel

in einer zähen Flüssigkeit  15–21

A. Iliev and D. Markushevich The Abel-Jacobi Map for a Cubic Threefold

and Periods of Fano Threefolds of Degree 14  23–47

N. Christopher Phillips A Classification Theorem

for Nuclear Purely Infinite Simple C^{∗}-Algebras  49–114

Eva Maria Feichtner and Günter M. Ziegler The Integral Cohomology Algebras of

Ordered Configuration Spaces of Spheres  115–139

Laurent Bonavero and Shigeharu Takayama Some Boundedness Results for

Fano-Like Moishezon Manifolds  141–150

Jinya Nakamura On the Milnor K-Groups of

Complete Discrete Valuation Fields  151–200

Elke Wilczok New Uncertainty Principles for the Continuous Gabor Transform

and the Continuous Wavelet Transform  201–226

Claudia Wulff Transitions from Relative Equilibria

to Relative Periodic Orbits  227–274

Jens Lieberum The Number of Independent Vassiliev Invariants in

the Homfly and Kauffman Polynomials  275–299

Heinz König On the Inner Daniell-Stone

and Riesz Representation Theorems  301–315

Erez Lapid and Jonathan Rogawski Stabilization of periods of Eisenstein series and

Bessel distributions on GL(3) relative to U(3)  317–350

Andrew Baker I_{n}-Local Johnson-Wilson Spectra

and their Hopf Algebroids  351–364

Hans-Georg Rück and Ulrich Tipp Heegner Points and L-Series of Automorphic

Cusp Forms of Drinfeld Type  365–444

J. F. Jardine

Motivic Symmetric Spectra  445–552

Ivan Kausz A Modular Compactification

of the General Linear Group  553–594

Ralph J. Bremigan

Pseudokähler Forms on Complex Lie Groups  595–612

Michael Brinkmeier Strongly Homotopy-Commutative

Monoids Revisited  613–624

Robert Lauter, Bertrand Monthubert, Victor Nistor Pseudodifferential Analysis

on Continuous Family Groupoids  625–655

M. V. Bondarko Local Leopoldt’s Problem for Rings of Integers in Abelian p-Extensions

of Complete Discrete Valuation Fields  657–693

Bernold Fiedler and Rolf M. Mantel

Crossover Collision of Scroll Wave Filaments  695–731